OpenStudy (anonymous):

Events A and B are overlapping. P(A) = 3/11 P(B) = 6/11 P(A and B) = 4/11 find p(a or b)

jimthompson5910 (jim_thompson5910):

Use the formula P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

13/11 ?

jimthompson5910 (jim_thompson5910):

no, 13/11 is larger than 1

jimthompson5910 (jim_thompson5910):

so a probability larger than 1 is always impossible

jimthompson5910 (jim_thompson5910):

P(A or B) = P(A) + P(B) - P(A and B) P(A or B) = 3/11 + 6/11 - 4/11 P(A or B) = (3 + 6 - 4)/11 P(A or B) = 5/11